# Trainable Dynamic Mask Sparse Attention

### v1

@@ 5.2 Variants Comparison
First, we present the comparison of the perplexity performance of different self-attention variants at various parameter scales in Figure 4. This experiment includes the baseline, sliding window attention driven by static mask structures, multi-head latent attention driven by low-rank decomposition approximations, native sparse attention  [^1] The implementation code for NSA is available at https://github.com/lucidrains/native-sparse-attention-pytorch.driven by hardware content adaptation, and our proposed Dynamic Mask Attention. These experiments were conducted on the SmolLMCorpus dataset, with model sizes ranging from 80M to 1.7B parameters, and the experimental configurations are detailed in Table 4. Our experimental results validate that Dynamic Mask Attention maintains the best performance across various scales. We speculate that this advantage primarily stems from DMA’s ability to adaptively focus on key information in the input sequence, effectively avoiding the lost in middle [34] problem.
@@
To further validate the ability of different attention variants in long-sequence information retrieval, we designed a more challenging variant of the multi-query associative recall task [2], which includes longer sequence lengths and smaller model dimensions. This task assesses the ability of language models to retrieve information within their context. Specifically, it provides key-value pairs to the autoregressive model, prompting the model to generate the correct value when displaying previously seen keys. To increase the difficulty of the task, we used 512 key-value pairs in the experiment. We employ sliding window attention, native sparse attention, and dynamic mask attention, all with a window size of 512. This approach replaces non-query/key/value parts with random tokens, forcing the model to locate relevant information precisely rather than relying on contextual clues. The experimental dataset comprises 250,000 training samples and 1,000 test samples, with all models trained for 100 epochs to ensure sufficient convergence. As shown in Figure 5, Dynamic Mask Attention performs excellently across various sequence lengths, indicating its ability to intelligently identify and focus on tokens relevant to the current state while ignoring irrelevant tokens.
Inference Speed.
Furthermore, as shown in Figure 6, DMA demonstrates significant advantages in inference speed for the multi-query associative recall task. Compared to the baseline MHA, DMA achieves substantial speed improvements across all tested sequence lengths. Notably, although DMA has a similar theoretical computational complexity to SWA, it requires additional sampling from the value state due to its unique dynamic mask mechanism, resulting in slightly slower speeds than SWA on shorter sequences. The inference speed of DMA is comparable to that of efficient variants of SWA, both having a computational complexity of $O(nwd_{h})$, where $n$ is the sequence length, $w$ is the window size, and $d_{h}$ is the head dimension. However, as the sequence length increases to 4096 and beyond, DMA’s efficiency catches up and even slightly surpasses that of SWA, indicating that the additional sampling overhead of DMA is significantly offset by its speed improvements on more extended sequences.
@@
Figure 5: Multi-Query Associative Recall. This is a more challenging version of the original multi-query associative recall task [2], which includes longer sequence lengths and smaller model dimensions. Dynamic Mask Attention maintains good performance in most cases.
Figure 6: Speed of Inference. The inference speed of these PyTorch implementations of variants when performing the multi-query associative recall task. Although both sliding window attention and dynamic mask attention have a computational complexity of $O(n\times w\times d_{h})$, dynamic mask attention requires additional sampling from the value state, making its speed slightly slower than sliding window attention on short sequences.
5.3 Implementation Comparison
Table 2: Forward Propagation Performance Comparison: SDPA vs DMA Implementations. Forward propagation performance comparison between SDPA baseline and DMA implementations in CUDA, Triton, and Flex. All times are in milliseconds.
Configuration SDPA DMA-CUDA DMA-Triton DMA-Flex Best Impl Best Speedup
Training Scenarios: batch=1, head=2, kv_head=1, d_h=128, w=2048
q_len=256, k_len=256 0.22 0.06 0.14 0.83 CUDA 3.9×
q_len=512, k_len=512 0.20 0.04 0.19 0.81 CUDA 5.0×
q_len=1024, k_len=1024 0.27 0.05 0.41 1.85 CUDA 5.5×
q_len=2048, k_len=2048 0.48 0.07 0.31 0.96 CUDA 6.6×
q_len=4096, k_len=4096 2.21 0.20 0.57 1.13 CUDA 10.9×
q_len=8192, k_len=8192 8.21 0.78 1.09 1.78 CUDA 10.6×
q_len=16384, k_len=16384 31.61 2.93 3.11 3.00 CUDA 10.8×
q_len=32768, k_len=32768 126.30 10.09 8.85 8.12 Flex 15.5×
Inference Scenarios: batch=1, head=2, k_len=v_head=1, q_len=1, d_h=128, w=2048
k_len=256 0.21 0.05 0.16 0.90 CUDA 4.3×
k_len=512 0.19 0.04 0.12 1.03 CUDA 5.1×
k_len=1024 0.19 0.04 0.12 1.01 CUDA 4.9×
k_len=2048 0.21 0.07 0.11 1.50 CUDA 3.2×
k_len=4096 0.22 0.06 0.13 1.01 CUDA 3.8×
k_len=8192 0.22 0.06 0.13 1.00 CUDA 3.9×
k_len=16384 0.31 0.06 0.12 0.89 CUDA 5.2×
k_len=32768 0.84 0.08 0.12 0.87 CUDA 11.2×
k_len=65536 1.63 0.10 0.13 1.01 CUDA 15.5×
Different Batch Sizes: head=2, kv_head=1, q_len=4096, k_len=4096, d_h=32, w=2048
batch=1 2.13 0.22 0.60 0.98 CUDA 9.9×
batch=2 4.15 0.40 0.49 1.24 CUDA 10.5×
batch=4 7.94 0.57 0.73 1.47 CUDA 13.9×
batch=8 15.74 1.52 1.22 1.57 Triton 12.9×
Different Head Numbers: batch=1, q_len=4096, k_len=4096, d_h=32, w=2048
head=1, kv_head=1 0.27 0.17 0.44 0.91 CUDA 1.6×
head=2, kv_head=1 2.12 0.20 0.59 0.92 CUDA 10.4×
head=4, kv_head=1 4.04 0.31 0.50 1.08 CUDA 12.9×
head=8, kv_head=2 8.00 0.46 0.69 1.25 CUDA 17.6×
Different Head Dimensions: batch=1, head=2, kv_head=1, q_len=4096, k_len=4096, w=2048
d_h=32 2.16 0.22 0.62 0.95 CUDA 9.9×
d_h=64 2.15 0.16 0.50 1.05 CUDA 13.9×
d_h=96 2.11 0.20 0.56 0.99 CUDA 10.5×
d_h=128 2.11 0.21 0.60 1.00 CUDA 10.1×
@@ 5.3 Implementation Comparison
To further analyze the performance of Dynamic Mask Attention in modern efficient operator frameworks, we provide a performance comparison of different kernel implementations on actual hardware. We compare the SDPA [43] baseline with DMA under three different kernel implementations: CUDA [38], Triton [49], and Flex [16]. All tests were conducted on an NVIDIA GeForce RTX 4090 GPU, with the average time taken from five runs after two warm-up iterations. As shown in Table 2, different implementations of DMA demonstrate significant performance advantages across various configurations. The table is divided into five sections: training scenarios, inference scenarios, tests with different batch sizes, tests with different numbers of attention heads, and tests with different head dimensions. The CUDA implementation performs best in most configurations, achieving over 10x speedup, especially on long sequences. Notably, as the sequence length increases, the speedup ratio shows an increasing trend, confirming DMA’s advantages in long-context processing scenarios.
@@ 5.4 Performance Comparison
Table 3: Downstream Task Zero-shot Evaluations. The best results for each size are in bold, and the second-best results are unlined. DMA outperforms MHA and NSA, as well as other advanced inference sparse methods, on most tasks.
Model Pile LAMBADA LAMBADA MMLU TriviaQA ARC PIQA HellaSwag OBQA WinoGrande LongBench
ppl $\downarrow$ ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ avg $\uparrow$
@@
MHA 48.65 15.22 44.3 35.4 9.4 53.4 72.9 56.1 37.0 57.3 14.2
H2O 15.38 44.2 34.8 7.4 53.3 72.8 55.6 36.6 56.9 8.7
InfLLM 48.96 15.23 44.2 35.1 8.0 53.1 72.4 55.8 36.6 56.8 9.2
Quest 49.68 15.43 43.9 35.1 7.6 53.1 72.6 56.1 36.8 57.2 9.6
DAM 49.72 15.89 44.5 34.6 8.9 52.1 72.3 56.2 36.3 56.0 10.4
Exact-Top 53.31 15.23 44.4 35.3 9.2 53.3 72.8 56.0 36.8 57.0 13.8
NSA 48.73 14.91 45.2 33.8 8.7 53.1 72.8 56.7 36.3 57.8 15.4
DMA (ours) 45.12 14.42 45.9 37.0 9.1 55.6 73.4 56.4 36.5 58.4 16.2
@@
MHA 19.40 40.4 36.8 13.2 56.8 73.2 56.8 38.0 58.6
H2O 19.14 38.9 35.7 10.2 56.6 73.2 56.4 37.8 58.1
InfLLM 19.13 41.3 35.9 11.7 56.7 73.3 56.1 38.0 57.7
Quest 19.22 40.9 36.1 10.9 56.2 73.2 55.8 37.9 58.2
DAM 19.47 41.2 35.2 13.3 55.1 71.0 54.4 38.0 57.2
Exact-Top 18.22 39.7 36.4 13.1 56.3 73.4 56.5 38.2 58.5
NSA 21.37 39.6 34.6 12.5 56.1 76.0 58.9 39.2 58.3
DMA (ours) 17.88 40.9 38.2 12.6 56.4 76.6 58.7 39.6 60.4
Downstream Benchmark Evaluations.
We used the Qwen3 1.7B [48] model structure as a baseline, making only modifications to the self-attention part for comparison. We first pre-trained the model on a high-quality dataset covering four domains: Web, TextBook, Code, and Math, with a total of 32 billion tokens and a sequence length of 2,048, thereby providing the model with basic language skills and general knowledge. Subsequently, we carefully selected 8B tokens packaged into sequences of length 8K. We conducted a second phase of pre-training by adjusting the RoPE base frequency from 10K to 100K [53], ensuring that the model could effectively handle longer inputs. Ultimately, we obtained three models: MHA, NSA, and DMA, and evaluated them on the following tasks: Pile [18], LLAMBADA [41], MMLU [25], TriviaQA [28], ARC [10], PIQA [6], HellaSwag [56], OBQA [37], Winogrande [45], and the English tasks of LongBench [3]. We also compared several advanced inference sparse methods, including H2O [59], infLLM [52], Quest [47], DAM [57], and Exact-Top, which first computes full attention scores using MHA and then performs sparsification based on that. The results are shown in Table 3. In both zero-shot and five-shot settings, DMA outperforms the baseline on most tasks, achieving excellent overall performance. This indicates that the sparse attention pre-training mechanism of Dynamic Mask Attention helps the model develop a special attention mechanism, as illustrated in Figure 9, which forces the model to focus on the most important information.
@@
We further conducted an apples-to-apples comparison between MHA, NSA, and DMA using the needle-in-a-haystack task [29] to evaluate the models’ ability to retrieve information accurately from long texts. In this synthetic retrieval task, a random and information-rich sentence is inserted into a lengthy document, and the model needs to retrieve the needle from the haystack to answer the question. As shown in Figure 7, as the context length increases, the advantage of DMA over NSA and MHA gradually expands. Notably, when the context length exceeds the pre-training sequence length, all three models exhibit a performance decline; however, the decrease in DMA’s performance is significantly smaller than that of NSA and MHA, demonstrating stronger extrapolation capabilities and more effective retrieval of information in unseen length ranges. We speculate that trainable sparse attention inherently possesses stronger sequence length extrapolation. This experimental result has dual significance: on one hand, it validates DMA’s intrinsic advantages in handling ultra-long documents, especially in practical application scenarios that require precise localization and extraction of key information; on the other hand, it reveals the structural advantages of DMA’s content-aware dynamic mask mechanism in maintaining long-distance dependency modeling capabilities, even when the sequence length exceeds the pre-training range, thus maintaining relatively stable performance. This extrapolation capability is of great value for practical applications that require processing long documents.
@@
Our comprehensive experimental results demonstrate the exceptional performance of Dynamic Mask Attention across various tasks and model scales. In scaling perplexity experiments, DMA consistently outperformed other attention variants across different parameter scales from 80M to 1.7B; in the multi-query associative recall task, DMA exhibited superior information retrieval capabilities and computational efficiency; in kernel implementations, DMA showed very high speedup ratios in various long-sequence application scenarios; in downstream benchmark evaluations, DMA models surpassed the original MHA and its various sparse variants on most tasks; in the needle-in-a-haystack task, DMA demonstrated significantly stronger length extrapolation capabilities. These results collectively validate the effectiveness of DMA as a sparse attention solution that simultaneously improves computational efficiency and model performance.

### v6

@@ 5.2 Variants Comparison
First, we present the comparison of the perplexity performance of different self-attention variants at various parameter scales in Figure 5. This experiment includes the baseline  [^1] The implementation code for MHA is available at https://github.com/Dao-AILab/flash-attention., sliding window attention  [^2] The implementation code for SWA is available at https://github.com/Dao-AILab/flash-attention. driven by static mask structures, multi-head latent attention  [^3] The implementation code for MLA is available at https://github.com/deepseek-ai/FlashMLA. driven by low-rank decomposition approximations, native sparse attention  [^4] The implementation code for NSA is available at https://github.com/lucidrains/native-sparse-attention-pytorch. driven by hardware content adaptation, and our proposed Dynamic Mask Attention. These experiments were conducted on the SmolLMCorpus dataset, with model sizes ranging from 80M to 1.7B parameters, and the experimental configurations are detailed in Table 4. Our experimental results validate that Dynamic Mask Attention maintains the best performance across various scales. We speculate that this advantage primarily stems from DMA’s ability to adaptively focus on key information in the input sequence, effectively avoiding the lost in middle [liu2023lostmiddlelanguagemodels] problem.
@@
To further validate the ability of different attention variants in long-sequence information retrieval, we designed a more challenging variant of the multi-query associative recall task [arora2024zoology], which includes longer sequence lengths and smaller model dimensions. This task assesses the ability of language models to retrieve information within their context. Specifically, it provides key-value pairs to the autoregressive model, prompting the model to generate the correct value when displaying previously seen keys. To increase the difficulty of the task, we used 512 key-value pairs in the experiment. We employ sliding window attention, native sparse attention, and dynamic mask attention, all with a window size of 512. This approach replaces non-query/key/value parts with random tokens, forcing the model to locate relevant information precisely rather than relying on contextual clues. The experimental dataset comprises 250,000 training samples and 1,000 test samples, with all models trained for 100 epochs to ensure sufficient convergence. As shown in Figure 6, Dynamic Mask Attention performs excellently across various sequence lengths, indicating its ability to intelligently identify and focus on tokens relevant to the current state while ignoring irrelevant tokens.
@@
Figure 6: Multi-Query Associative Recall. This is a more challenging version of the original multi-query associative recall task [arora2024zoology], which includes longer sequence lengths and smaller model dimensions. Dynamic Mask Attention maintains good performance in most cases.
Figure 7: Kernal Performance. A performance comparison of efficient implementation kernels of different attention variants on an A100 GPU. Our DMA achieves significant acceleration in forward propagation, decoding, and backward propagation, maintaining the same efficiency level as SWA.
@@
To analyze the performance of DMA within modern efficient operator frameworks, we benchmarked the forward, decoding, and backward performance of MHA, SWA, MLA, NSA, and DMA on an A100-SXM4-80GB GPU. The results represent the average of 1,000 runs after three warm-up iterations; specific configurations and implementations can be found in Table 5. As shown in Figure 7, for the Forward pass, compared to MHA, DMA achieves speedups of approximately $26.1\times$, $10.2\times$, and $21.5\times$ at token lengths of $8192$, $16384$, and $32768$, respectively. For the Decode phase, speedups against MHA at key lengths of $65536$, $131072$, $262144$, and $524288$ are approximately $49.6\times$, $92.7\times$, and $171.1\times$, respectively. For the Backward pass, speedups against MHA at lengths of $8192$, $16384$, and $32768$ are approximately $2.5\times$, $4.4\times$, and $7.9\times$, respectively. DMA avoids large-scale redundant score and softmax backpropagation overhead by explicitly skipping masked blocks, thus providing strong acceleration capabilities across multiple critical stages.
@@ 5.3 Performance Comparison
Table 2: Downstream Task Evaluations for Base Model. The best results for each size are in bold, and the second-best results are unlined. DMA outperforms MHA and NSA, as well as other advanced inference sparse methods, on most tasks.
Model LAMBADA LAMBADA MMLU TriviaQA ARC PIQA HellaSwag OBQA WinoGrande Avg
ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
@@
MHA 15.22 44.3 35.4 9.4 53.4 72.9 56.1 37.0 57.3 45.7
H2O 15.38 44.2 34.8 7.4 53.3 72.8 55.6 36.6 56.9 45.2
InfLLM 15.23 44.2 35.1 8.0 53.1 72.4 55.8 36.6 56.8 45.2
Quest 15.43 43.9 35.1 7.6 53.1 72.6 56.1 36.8 57.2 45.3
DAM 15.89 44.5 34.6 8.9 52.1 72.3 56.2 36.3 56.0 45.1
Exact-Top 15.23 44.4 35.3 9.2 53.3 72.8 56.0 36.8 57.0 45.6
NSA 14.91 45.2 33.8 8.7 53.1 72.8 56.7 36.3 57.8 45.5
DMA (ours) 14.42 45.9 37.0 9.1 55.6 73.4 56.4 36.5 58.4 46.5
@@
MHA 19.40 40.4 36.8 13.2 56.8 73.2 56.8 38.0 58.6 46.7
H2O 19.14 38.9 35.7 10.2 56.6 73.2 56.4 37.8 58.1 45.8
InfLLM 19.13 41.3 35.9 11.7 56.7 73.3 56.1 38.0 57.7 46.3
Quest 19.22 40.9 36.1 10.9 56.2 73.2 55.8 37.9 58.2 46.1
DAM 19.47 41.2 35.2 13.3 55.1 71.0 54.4 38.0 57.2 45.6
Exact-Top 18.22 39.7 36.4 13.1 56.3 73.4 56.5 38.2 58.5 46.5
NSA 21.37 39.6 34.6 12.5 56.1 76.0 58.9 39.2 58.3 46.9
DMA (ours) 17.88 40.9 38.2 12.6 56.4 76.6 58.7 39.6 60.4 47.9
Table 3: Downstream Evaluations for Finetuned Model. The best results are in bold, and the second-best results are underlined. For models supervisedly fine-tuned at a 16K sequence length, DMA outperforms other methods in most tasks.
Method MMLU BBH GSM8K MATH MBPP ARC PIQA LongBench RULER Avg
acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
MHA 46.4 37.7 46.3 11.7 40.0 59.8 76.2 30.2 60.6 39.0
H2O 42.4 34.5 44.8 10.5 38.5 57.6 74.1 28.5 47.8 35.3
InfLLM 44.2 36.0 45.5 11.0 39.2 58.9 75.6 29.1 55.3 37.2
Quest 44.0 34.8 45.0 10.8 38.8 56.2 73.7 28.9 50.7 36.1
DAM 45.0 36.5 46.0 11.5 37.2 56.9 72.4 26.4 49.2 36.0
Exact-Top 45.1 37.2 45.1 11.5 38.4 57.3 75.4 28.3 44.8 35.7
NSA 43.8 38.4 46.2 11.6 40.5 59.1 75.7 30.2 59.6 38.6
DMA (ours) 46.2 38.2 46.8 11.6 40.6 59.6 76.6 30.7 60.5 39.2
Downstream Benchmark Evaluations for Base Model.
We used the Qwen3 1.7B [qwen32025] model structure as a baseline, making only modifications to the self-attention part for comparison. We first pre-trained the model on a high-quality dataset covering four domains: Web, TextBook, Code, and Math, with a total of 32 billion tokens and a sequence length of 2,048, thereby providing the model with basic language skills and general knowledge. Subsequently, we carefully selected 8B tokens packaged into sequences of length 8K. We conducted a second phase of pre-training by adjusting the RoPE base frequency from 10K to 100K [xiong2023effectivelongcontextscalingfoundation], ensuring that the model could effectively handle longer inputs. Ultimately, we obtained three models: MHA, NSA, and DMA, and evaluated them on the following tasks: LLAMBADA [paperno2016lambada], MMLU [hendrycks2021measuring], TriviaQA [joshi2017triviaqa], ARC [clark2018think], PIQA [bisk2020piqa], HellaSwag [zellers2019hellaswag], OBQA [mihaylov2018can], Winogrande [sakaguchi2021winogrande], and the English tasks of LongBench [bai2023longbench]. We also compared several advanced inference sparse methods, including H2O [zhang2023h2o], infLLM [xiao2024infllm], Quest [tang2024quest], DAM [zhang2025damdynamicattentionmask], and Exact-Top, which first computes full attention scores using MHA and then performs sparsification based on that. The results are shown in Table 2. In both zero-shot and five-shot settings, DMA outperforms the baseline on most tasks, achieving excellent overall performance.
Downstream Benchmark Evaluations for Finetuned Model.
We further fine-tuned all models at a 16K sequence length with an adjusted RoPE base frequency to 400K, to further enhance the models’ long-context generalization capabilities. Ultimately, we obtained three fine-tuned models: MHA, NSA, and DMA, and evaluated them on the following tasks: MMLU [hendrycks2021measuring], BBH [suzgun2023challenging], GSM8K [cobbe2021training], MATH [hendrycks2020measuring], MBPP [austin2021program], LongBench [bai2023longbench], and RULER [hsieh2024ruler]. We used the same advanced inference sparse methods, which also compute full attention scores using MHA and then perform sparsification based on that. The results are shown in Table 3. DMA achieved the best average score, leading in GSM8K, MBPP, and LongBench, while remaining highly competitive in MMLU and RULER. NSA ranked first in BBH, with DMA closely following, while the full-attention MHA performed best in RULER but lagged behind DMA in average score. These results indicate that DMA’s content-aware sparse mask effectively transfers even under longer context fine-tuning.
@@
We further conducted an apples-to-apples comparison between MHA, NSA, and DMA using the needle-in-a-haystack task [gkamradt2023needleinahaystack] to evaluate the models’ ability to retrieve information accurately from long texts. In this synthetic retrieval task, a random and information-rich sentence is inserted into a lengthy document, and the model needs to retrieve the needle from the haystack to answer the question. As shown in Figure 8, as the context length increases, the advantage of DMA over NSA and MHA gradually expands. Notably, when the context length exceeds the pre-training sequence length, all three models exhibit a performance decline; however, the decrease in DMA’s performance is significantly smaller than that of NSA and MHA, demonstrating stronger extrapolation capabilities and more effective retrieval of information in unseen length ranges. We speculate that trainable sparse attention inherently possesses stronger sequence length extrapolation. This experimental result has dual significance: on one hand, it validates DMA’s intrinsic advantages in handling ultra-long documents, especially in practical application scenarios that require precise localization and extraction of key information; on the other hand, it reveals the structural advantages of DMA’s content-aware dynamic mask mechanism in maintaining long-distance dependency modeling capabilities, even when the sequence length exceeds the pre-training range, thus maintaining relatively stable performance. This extrapolation capability is of great value for practical applications that require processing long documents.
@@
Our comprehensive experimental results demonstrate the exceptional performance of Dynamic Mask Attention across various tasks and model scales. In scaling perplexity experiments, DMA consistently outperformed other attention variants across different parameter scales from 80M to 1.7B; in the multi-query associative recall task, DMA exhibited superior information retrieval capabilities and efficiency; in kernel implementations, DMA showed extremely high speedup ratios in various long-sequence application scenarios; in downstream benchmark evaluations, DMA models outperformed the original MHA and its various sparse variants on most tasks; in the needle-in-a-haystack task, DMA demonstrated significantly stronger length extrapolation capabilities. These results collectively validate the effectiveness of DMA as a sparse attention solution that simultaneously enhances computational efficiency and model performance.
Figure 9: Weights Heatmaps of DMA. The heatmaps show the attention weights of each head in the Dynamic Mask Attention mechanism, indicating which tokens each head focuses on. Full heatmaps can be found in Appendix C.
Figure 10: Adaptive Filtering of DMA. Top: the noisy token signal retains its underlying low-frequency structure while the adaptive mask focuses on informative tokens compared with uniform sampling. Bottom: the learned content-aware scores and the resulting mask illustrate how DMA allocates compute to relevant regions while skipping redundant context.
6 Analysis
In this section, we analyze Dynamic Mask Attention, highlighting its distinct advantages in handling long-range dependencies and providing dynamic context awareness.
6.1 Head Specialization
As shown in Figure 9 11, our analysis of the attention patterns learned by the model reveals how DMA creates content-aware sparse structures that adapt to different contextual needs. Unlike the uniform patterns of traditional attention mechanisms, each DMA attention head develops a unique sparse pattern: some heads focus on the most recent tokens to capture local context, while others attend to specific distant positions for long-range dependencies, and additional heads maintain broader context awareness for global understanding. This diversity allows the model to capture various types of dependencies simultaneously while maintaining computational efficiency, making efficient use of each subspace.
Local Context Heads. For example, Head 0, Head 1, Head 8, Head 10, and Head 11 tend to focus on the most recent tokens, forming a local band attention pattern. These heads are primarily responsible for capturing syntactic structures, phrase-level semantics, and local dependencies, which are particularly important for tasks requiring precise local context handling.
Range Dependency Heads. For example, Head 2, Head 3, Head 4, Head 5, and Head 14 demonstrate the ability to attend to specific distant tokens. These heads are specialized in capturing long-range semantic associations, such as resolving coreference issues or tracking complex storylines in lengthy documents. They can skip over large amounts of intermediate information and directly connect distant but semantically related parts, which is crucial for deep reasoning and contextual understanding.
Global Context Heads. For example, Head 6, Head 7, Head 9, Head 12, Head 13, and Head 15 exhibit a sparser but broader attention distribution, sampling key information from the entire sequence to form an overall perception of the global context. These heads function similarly to summarizers, responsible for integrating information from different parts to create a coherent global representation. This capability is crucial for tasks that require a comprehensive understanding of the entire input to make accurate predictions.
Dynamic Adaptability. The most significant advantage of DMA lies in its dynamism. These attention patterns are not static; they are dynamically generated based on the input content. This means the model can adjust its attention strategy in real-time, activating the most appropriate combination of heads when processing different tasks or text types. For example, when processing code, it might rely more on long-range dependency heads to track variable definitions and usages, whereas in a conversation, it might focus more on local context heads to understand the current exchange. This content-aware adaptability is the core advantage of DMA over static sparse attention methods.
This naturally occurring specialization is a direct result of the content-aware mask mechanism, enabling the model to effectively handle various types of dependencies while maintaining computational efficiency, achieving effective integration of multi-scale information. This hierarchical integration mechanism can effectively handle multi-level semantic structures in complex texts. It is worth noting that head specialization may also occur in traditional MHA, but the specialization patterns in DMA are more pronounced and functionally clearer, which may be a key reason for its superior performance across various tasks.
6.2 Adaptive Filtering
From the perspective of modern signal processing, Trainable Dynamic Mask Sparse Attention essentially performs dynamic downsampling of the input sequence through learnable adaptive filters, i.e., masks, retaining only key information components. This allows for efficient extraction of low-frequency dependencies in long-distance signals, such as text, while suppressing noise redundancy. The core logic is to treat long texts as noisy low-frequency signals, where the mask acts as an adaptive filter, and the dynamically selected retained tokens are equivalent to intelligent downsampling based on signal relevance.
Learnable Content-aware Filters. Unlike traditional sparse methods that rely on fixed patterns, DMA’s mask is dynamically generated based on the input content, making it a content-aware adaptive filter. This filter learns to identify and amplify key components in the signal, i.e., important tokens, while attenuating or completely filtering out noise, i.e., irrelevant tokens. This mechanism ensures that computational resources are precisely allocated to the most critical information for the current task, effectively avoiding information loss due to excessively long contexts in needle-in-a-haystack problems.
Multi-scale Signal Decomposition. The different attention heads in DMA learn various sparse patterns, which can be viewed as a set of parallel adaptive filters, each responsible for capturing different scales or types of signal features. This multi-scale decomposition allows the model to build a comprehensive and hierarchical understanding of the input signal with extremely high efficiency.
Recasting sparse attention as an adaptive filtering problem, DMA offers a new perspective for understanding and optimizing long text processing. It achieves intelligent filtering of information through content awareness and multi-scale decomposition, ensuring that the model learns the optimal sparse strategies.